# Noise-induced resonant acceleration of a charge in an intermittent magnetic field: an exact solution for ergodic and non-ergodic fluctuations

### v1

@@ II Trajectory approach: general exact solution
Figure 2: Diffusion of a charged particle in a fluctuating dichotomous magnetic field $B(t)$ in the Poissonian case. The green line is $\langle r^{2}\rangle$ from solution (27), blue and red line are $\langle x\rangle$, $\langle y\rangle$ from solution (S18). Black diamonds, triangles and circles are the respective simulated results, averaged over 50k runs, which show perfect agreement. Parameters are $v_{0}=1,\gamma=20,\,\omega_{0}=1,\,\omega_{1}=0.7$ and initial conditions $x(0)=y(0)=\dot{x}(0)=0$ and $\dot{y}(0)=v_{0}$ corresponding to ${\bf A}^{\pm}_{0}=\frac{1}{2}(\frac{1}{\omega_{\pm}},-\frac{1}{|\omega_{\pm}|})$. Inset: corresponding solutions from the density approach Eqs.(S32), (S31). Setting $\omega_{0}$=0 in (S32) and (27) (brown line) coincides with prediction of Eq. (34) and with simulation (triangles), showing confinement at any time.
@@
$=\sum_{i,j,k_{1},k_{2},k_{3},k_{4}}{\bf A^{\dagger}}_{0}(k_{2})M^{\dagger}_{k_{2},k_{1}}(\tau_{1})M^{\dagger}_{k_{1},i}(\tau_{2})\sigma_{ij}(t-\tau_{2})$
@@
$|{\bf u}_{2}(t)|^{2}={\bf\Sigma}(t-\tau_{2},\tau_{2})\cdot\tilde{M}(\tau_{1}){\bf\tilde{A}_{0}}$ (21)
@@
$\Sigma(t-\tau_{2},\tau_{2})=M^{\dagger}(\tau_{2})\sigma(t-\tau_{2})M(\tau_{2})$ (22)
@@
$\int_{\tau_{1}}^{t}d\tau_{2}\int_{0}^{t}d\tau_{1}\Psi(t-\tau_{2})\psi(\tau_{2}){\bf\Sigma}(t-\tau_{2},\tau_{2})\cdot\tilde{M}(\tau_{1})\psi(\tau_{1}){\bf\tilde{A}_{0}}$
@@
$\langle r^{2}(t)\rangle=\sum_{n=0}^{\infty}\int_{\tau_{n-1}}^{t}d\tau_{n}\int_{\tau_{n-2}}^{t}d\tau_{n-1}\dots\int_{0}^{t}d\tau_{1}\Psi(t-\tau_{n})\;\;\;\;\;\;$
$\psi(\tau_{n}){\bf\Sigma}(t-\tau_{n},\tau_{n})\cdot\tilde{M}(\tau_{n-1})\psi(\tau_{n-1})\cdots\tilde{M}(\tau_{1})\psi(\tau_{1}){\bf\tilde{A}_{0}}$
@@
where $\tilde{M}_{\psi_{\pm}}(s)$ is the Laplace transform $\mathcal{L}_{s}[\psi(t)\tilde{M}_{\pm}(t)]$ and ${\bf\tilde{\Sigma}_{\pm}}(s)=\mathcal{L}_{s}\left[\Psi(t-\tau)\psi(\tau){{\bf\Sigma}_{\pm}}(t-\tau,\tau)\right]$. Eqs. (25) can therefore be expressed as a function only of the Laplace Transform of $\psi(\tau)$, and then used to evaluate $\langle r^{2}(t)\rangle$ via Eq. (16). For Poissonian fluctuations of rate $\gamma$ this leads to
@@
Slightly simpler is the derivation of $\langle z\rangle$ and therefore of $\langle x,y\rangle$, starting from trajectories of $u$ (see Appendix A2 for details). Figure 2 shows $\langle r^{2}\rangle$$\langle x\rangle$ and $\langle y\rangle$ so evaluated, which coincide with the result from the density approach (inset) and with the numerical simulation.
@@ II.3 Slow fluctuations and ergodicity breaking
Figure 3 compares for both cases the analytical prediction obtained by inverting Eq. (25) with the numerical simulations. Although the closed-form expressions in both cases are algebraically involved, their Laplace transform (see Eq. (S21), Eq. (S24) and Fig. 6 in the Appendix) exhibit a single positive pole that can be computed exactly. This pole determines the exponential growth rate of $\langle r^{2}(t)\rangle$, corresponding to the slopes of the black lines in Fig. 3.
@@ III The periodic case
For most values of $T$ (see Fig. 4 and Appendix A1 for details), the particle remains confined. However, for specific values of TT—the so-called ’resonance bands’—the switching frequency of the field resonates with the particle’s natural orbital motion, resulting in sustained acceleration. In the stochastic case, random fluctuations are always able to excite these resonance bands, as long as their spectrum is non-zero for frequencies corresponding to the time scales falling within the bands, thus providing a clear explanation for the robustness of the ’noise-induced resonant acceleration’ mechanism. Once a finite background field is present ($\omega_{0}>0$), confinement is broken and the particle is driven into an accelerated diffusion regime.
@@ IV Conclusions
V TRAJECTORY APPROACH: Case of periodic dichotomous fluctuations
@@
$|u_{2n}(t=2nT)|^{2}={\bf A_{0}}^{\dagger}\left(M_{\pm}^{\dagger}(T)M_{\mp}^{\dagger}(T)\right)^{n}\cdot\begin{bmatrix}1&1\\ 1&1\end{bmatrix}\cdot\left(M_{\pm}(T)M_{\mp}(T)\right)^{n}{\bf A_{0}}$ (S1)
@@
$\lambda_{\pm}=\frac{1}{\omega_{0}^{2}-\omega_{1}^{2}}\Biggl(\omega_{0}^{2}\cos[\omega_{0}T]-\omega_{1}^{2}\cos[\omega_{1}T]+$ (S2)
@@
Figure S1: Examples of trajectories generated with $\omega_{0}=1$ and $\omega_{1}=0.7$ at t=10000 for the periodic case with period $T=1$ (left, showing confinement) and $T=3$ (right, showing resonance).
For $\omega_{0}=0$, for instance, $\lambda_{\pm}=e^{\pm i\omega_{1}T}$ and their modulus is one, meaning confined motion. It is easy to check from (S2) that this condition of confinement (i.e. $|\lambda_{\pm}|=1$), applies to every value of $\omega_{0},\omega_{1}$ and $T$ for which the expression within square root is negative. When on the contrary, this is positive, one of the eigenvalues necessarily has modulus larger than 1, which leads for $n$ large to an exponentially diverging value for $\langle r^{2}\rangle$ which corresponds to resonance. Example of such conditions are shown in Figure S1, while an example of the resulting ’resonance bands’ (occurring for values of $\omega_{0},\omega_{1}$ and $T$ such that the expression within square root in (S2) is positive) for different values of the period T are shown in Figure 4 in main text.
V.1 TRAJECTORY APPROACH: stochastic fluctuations
V.1.1 Evaluation of $\langle z(t)\rangle$
@@
$u^{\pm}(t)=A^{\pm}_{n}e^{i\omega_{\pm}(t-t_{n})}+B^{\pm}_{n}e^{-i\omega_{\pm}(t-t_{n})}\;\;\;\text{for}\;\;\;t_{n}<t<t_{n+1}\\$ (S3)
@@
$u^{\pm}(t)=\left(e^{i\omega_{\pm}(t-t_{n})},e^{-i\omega_{\pm}(t-t_{n})}\right)\cdot{\bf A^{\pm}_{n}}$ (S4)
@@
$u^{\pm}(t)=\left(e^{i\omega_{\pm}(t-t_{n})},e^{-i\omega_{\pm}(t-t_{n})}\right)\cdot M_{\pm}(\tau_{n})\dots M_{\pm}(\tau_{1}){\bf A^{\pm}_{0}}$ (S5)
@@
$z(t)=e^{-i\left(\omega_{+}(t-t_{n})+\omega_{-}(t_{n}-t_{n-1})\cdots+\omega_{+}t_{1}\right)}u(t)$ (S6)
and therefore Eq. (S3) converts for $z$ into:
$z^{\pm}(t)=\left(1,e^{-2i\omega_{\pm}(t-t_{n})}\right)\cdot e^{-i\omega_{\pm}\tau_{n}}M_{\pm}(\tau_{n})\dots e^{-i\omega_{\pm}\tau_{1}}M_{\pm}(\tau_{1}){\bf A^{\pm}_{0}}$ (S7)
@@
$\langle z^{\pm}(t)\rangle=\sum_{n=0}^{\infty}\int_{0}^{t}d\tau\left[\left(\Psi_{+}(t-\tau),\Psi^{*}_{+}(t-\tau)\right)\cdot{\bf\Phi}_{2n}^{\pm\pm}(\tau)+\left(\Psi_{-}(t-\tau),\Psi^{*}_{-}(t-\tau)\right)\cdot{\bf\Phi}_{2n+1}^{\mp\pm}(\tau)\right]$ (S9)
@@
${\bf\Phi}_{2n}^{\pm\pm}(\tau)=\int_{0}^{\tau}d\tau_{2n}e^{-i\omega_{\mp}\tau_{2n}}\psi_{\mp}(\tau_{2n})M_{\mp}(\tau_{2n})\cdots\int_{0}^{\tau}d\tau_{1}e^{-i\omega_{\pm}\tau_{1}}\psi_{\pm}(\tau_{1})M_{\pm}(\tau_{1}){\bf A}^{\pm}_{0}$ (S10)
is the convolution of $2n$ functions $e^{-i\omega_{\pm}\tau}\psi_{\pm}(\tau)M_{\pm}(\tau)$ in correspondence of $2n$ magnetic field switches in the interval $(0,\tau)$ starting with $\pm$ and ending with same sign $\pm$. We are therefore summing over all the possible realisations, or trajectories, of the fluctuations of the field. The sum in Eq. (S9) can be evaluated in Laplace transform since, using the properties of convolution, from (S10) it follows:
$\hat{\Phi}_{2n}^{\pm\pm}(s)=\left(M_{\psi_{\mp}}(s+i\omega_{\mp})\right)^{n}\left(M_{\psi_{\pm}}(s+i\omega_{\pm})\right)^{n}{\bf A}^{\pm}_{0}\;\;,\;\;\;\hat{\Phi}_{2n+1}^{\pm\pm}(s)=M_{\psi_{\pm}}(s+i\omega_{\pm})\hat{\Phi}_{2n}^{\pm\pm}(s)$ (S11)
where $M_{\psi_{\pm}}(s)=\mathcal{L}_{s}[\psi_{\pm}(\tau)M_{\pm}(\tau)]$. This leads to a geometric series in (S9) in Laplace domain that can be evaluated exactly, from the Laplace Transform of (S9):
@@
$\langle\hat{z}(s)\rangle=\frac{\langle\hat{z}^{+}(s)\rangle+\langle\hat{z}^{-}(s)\rangle}{2}$ (S13)
@@
$\langle\hat{z}(s)\rangle=\frac{\omega_{1}(s+\gamma/2)}{s^{4}+2s^{3}(i\omega_{0}+\gamma)+s^{2}(\omega_{1}^{2}-\omega_{0}^{2}+3\omega_{1}\gamma+\gamma^{2})+s\gamma(\omega_{1}^{2}-\omega_{0}^{2}+i\omega_{0}\gamma)+\omega_{1}^{2}\gamma^{2}/4}$ (S14)
@@
$\langle z(t)\rangle=\omega_{1}\sum_{i=1}^{4}\frac{e^{\lambda_{i}t}(2\lambda_{i}+\gamma)}{(\lambda_{i}-\lambda_{j\neq i})(\lambda_{i}-\lambda_{k\neq j\neq i})(\lambda_{i}-\lambda_{m\neq k\neq j\neq i})}$ (S15)
@@
$x^{4}+2(\gamma+i\omega_{0})x^{3}+(\omega_{1}^{2}-\omega_{0}^{2}+3i\omega_{0}\gamma+\gamma^{2})x^{2}+\gamma(\omega_{1}^{2}-\omega_{0}^{2}+i\omega_{0}\gamma)x+\omega_{1}^{2}\gamma^{2}/4=0$ (S16)
@@
$\langle\hat{z}(s)\rangle=\frac{i(s+\gamma)(s+i\omega_{1}+\gamma)}{(s^{4}+2s^{3}(i\omega_{0}+\gamma)+s^{2}(\omega_{1}^{2}-\omega_{0}^{2}+3\omega_{0}\gamma+\gamma^{2})+s\gamma(\omega_{1}^{2}-\omega_{0}^{2}+i\omega_{0}\gamma)+\omega_{1}^{2}\gamma^{2}/4)}$ (S17)
@@
$\langle z(t)\rangle=i\sum_{i=1}^{4}\frac{e^{\lambda_{i}t}(\lambda_{i}^{2}+\lambda_{i}(2i\gamma-\omega_{0})+i\gamma^{2}-\omega_{0}\gamma)}{(\lambda_{i}-\lambda_{j\neq i})(\lambda_{i}-\lambda_{k\neq j\neq i})(\lambda_{i}-\lambda_{m\neq k\neq j\neq i})}$ (S18)
with $\lambda_{i}$ the four roots of the same equation (S16).
V.1.2 Evaluation of $\langle r^{2}(t)\rangle$: the case of non-ergodic and ergodic power law fluctuations

### v3

@@ II Trajectory approach: general exact solution
Figure 2: Diffusion of a charged particle in a fluctuating dichotomous magnetic field $B(t)$ in the Poissonian case. The green line is $\langle r^{2}\rangle$ from solution (27), blue and red line are $\langle x\rangle$, $\langle y\rangle$ from solution (A18). Black diamonds, triangles and circles are the respective simulated results, averaged over 50k runs, which show perfect agreement. Parameters are $v_{0}=1,\gamma=20,\,\omega_{0}=1,\,\omega_{1}=0.7$ and initial conditions $x(0)=y(0)=\dot{x}(0)=0$ and $\dot{y}(0)=v_{0}$ corresponding to ${\bf A}^{\pm}_{0}=\frac{1}{2}(\frac{1}{\omega_{\pm}},-\frac{1}{|\omega_{\pm}|})$. Inset: corresponding solutions from the density approach Eqs.(A32), (A31). Setting $\omega_{0}$=0 in (A32) and (27) (brown line) coincides with prediction of Eq. (34) and with simulation (triangles), showing confinement at any time.
@@
$=\sum_{i,j,k_{1},k_{2},k_{3},k_{4}}{\bf A^{\dagger}}_{0}(k_{2})M^{\dagger}_{k_{2},k_{1}}(\tau_{1})M^{\dagger}_{k_{1},i}(\tau_{2})\sigma_{ij}(t-t_{2})$
@@
$|{\bf u}_{2}(t)|^{2}={\bf\Sigma}(t-t_{2},\tau_{2})\cdot\tilde{M}(\tau_{1}){\bf\tilde{A}_{0}}$ (21)
@@
$\Sigma(t-t_{2},\tau_{2})=M^{\dagger}(\tau_{2})\sigma(t-t_{2})M(\tau_{2})$ (22)
@@
$\int_{t_{1}}^{t}dt_{2}\int_{0}^{t}dt_{1}\Psi(t-t_{2})\psi(\tau_{2}){\bf\Sigma}(t-t_{2},\tau_{2})\cdot\tilde{M}(\tau_{1})\psi(\tau_{1}){\bf\tilde{A}_{0}}$
@@
$\langle r^{2}(t)\rangle=\sum_{n=0}^{\infty}\int_{t_{n-1}}^{t}dt_{n}\int_{t_{n-2}}^{t}dt_{n-1}\dots\int_{0}^{t}dt_{1}\Psi(t-t_{n})\;\;\;\;\;\;$
$\psi(\tau_{n}){\bf\Sigma}(t-t_{n},\tau_{n})\cdot\tilde{M}(\tau_{n-1})\psi(\tau_{n-1})\cdots\tilde{M}(\tau_{1})\psi(\tau_{1}){\bf\tilde{A}_{0}}$
@@
where $\tilde{M}_{\psi_{\pm}}(s)$ is the Laplace transform $\mathcal{L}_{s}[\psi(t)\tilde{M}_{\pm}(t)]$ and ${\bf\tilde{\Sigma}_{\pm}}(s)=\mathcal{L}_{s}\left[\int_{0}^{t}d\tau\Psi(t-\tau)\psi(\tau){{\bf\Sigma}_{\pm}}(t-\tau,\tau)\right]$. Eqs. (25) can therefore be expressed as a function only of the Laplace Transform of $\psi(\tau)$, and then used to evaluate $\langle r^{2}(t)\rangle$ via Eq. (16). For Poissonian fluctuations of rate $\gamma$ this leads to
@@
Slightly simpler is the derivation of $\langle z\rangle$ and therefore of $\langle x,y\rangle$, starting from trajectories of $u$ (see Appendix  S2 for details). Figure 2 shows $\langle r^{2}\rangle$$\langle x\rangle$ and $\langle y\rangle$ so evaluated, which coincide with the result from the density approach (inset) and with the numerical simulation.
@@ II.3 Slow fluctuations and ergodicity breaking
Figure 3 compares for both cases the analytical prediction obtained by inverting Eq. (25) with the numerical simulations. Although the closed-form expressions in both cases are algebraically involved, their Laplace transform (see Eq. (A21), Eq. (A24) and Fig. A2 in the Appendix) exhibit a single positive pole that can be computed exactly. This pole determines the exponential growth rate of $\langle r^{2}(t)\rangle$, corresponding to the slopes of the black lines in Fig. 3.
@@ III The periodic case
For most values of $T$ (see Fig. 4 and Appendix S1 for details), the particle remains confined. However, for specific values of TT—the so-called ’resonance bands’—the switching frequency of the field resonates with the particle’s natural orbital motion, resulting in sustained acceleration. In the stochastic case, random fluctuations are always able to excite these resonance bands, as long as their spectrum is non-zero for frequencies corresponding to the time scales falling within the bands, thus providing a clear explanation for the robustness of the ’noise-induced resonant acceleration’ mechanism. Once a finite background field is present ($\omega_{0}>0$), confinement is broken and the particle is driven into an accelerated diffusion regime.
@@ IV Conclusions
V S1: TRAJECTORY APPROACH: Case of periodic dichotomous fluctuations
@@
$|u_{2n}(t=2nT)|^{2}={\bf A_{0}}^{\dagger}\left(M_{\pm}^{\dagger}(T)M_{\mp}^{\dagger}(T)\right)^{n}\cdot\begin{bmatrix}1&1\\ 1&1\end{bmatrix}\cdot\left(M_{\pm}(T)M_{\mp}(T)\right)^{n}{\bf A_{0}}$ (A1)
@@
$\lambda_{\pm}=\frac{1}{\omega_{0}^{2}-\omega_{1}^{2}}\Biggl(\omega_{0}^{2}\cos[\omega_{0}T]-\omega_{1}^{2}\cos[\omega_{1}T]+$ (A2)
@@
Figure A1: Examples of trajectories generated with $\omega_{0}=1$ and $\omega_{1}=0.7$ at t=10000 for the periodic case with period $T=1$ (left, showing confinement) and $T=3$ (right, showing resonance).
For $\omega_{0}=0$, for instance, $\lambda_{\pm}=e^{\pm i\omega_{1}T}$ and their modulus is one, meaning confined motion. It is easy to check from (A2) that this condition of confinement (i.e. $|\lambda_{\pm}|=1$), applies to every value of $\omega_{0},\omega_{1}$ and $T$ for which the expression within square root is negative. When on the contrary, this is positive, one of the eigenvalues necessarily has modulus larger than 1, which leads for $n$ large to an exponentially diverging value for $\langle r^{2}\rangle$ which corresponds to resonance. Example of such conditions are shown in Figure A1, while an example of the resulting ’resonance bands’ (occurring for values of $\omega_{0},\omega_{1}$ and $T$ such that the expression within square root in (A2) is positive) for different values of the period T are shown in Figure 4 in main text.
VI S2: TRAJECTORY APPROACH: stochastic fluctuations
VI.0.1 Evaluation of $\langle z(t)\rangle$
@@
$u^{\pm}(t)=A^{\pm}_{n}e^{i\omega_{\pm}(t-t_{n})}+B^{\pm}_{n}e^{-i\omega_{\pm}(t-t_{n})}\;\;\;\text{for}\;\;\;t_{n}<t<t_{n+1}\\$ (A3)
@@
$u^{\pm}(t)=\left(e^{i\omega_{\pm}(t-t_{n})},e^{-i\omega_{\pm}(t-t_{n})}\right)\cdot{\bf A^{\pm}_{n}}$ (A4)
@@
$u^{\pm}(t)=\left(e^{i\omega_{\pm}(t-t_{n})},e^{-i\omega_{\pm}(t-t_{n})}\right)\cdot M_{\pm}(\tau_{n})\dots M_{\pm}(\tau_{1}){\bf A^{\pm}_{0}}$ (A5)
@@
$z(t)=e^{-i\left(\omega_{+}(t-t_{n})+\omega_{-}(t_{n}-t_{n-1})\cdots+\omega_{+}t_{1}\right)}u(t)$ (A6)
and therefore Eq. (A3) converts for $z$ into:
$z^{\pm}(t)=\left(1,e^{-2i\omega_{\pm}(t-t_{n})}\right)\cdot e^{-i\omega_{\pm}\tau_{n}}M_{\pm}(\tau_{n})\dots e^{-i\omega_{\pm}\tau_{1}}M_{\pm}(\tau_{1}){\bf A^{\pm}_{0}}$ (A7)
@@
$\langle z^{\pm}(t)\rangle=\sum_{n=0}^{\infty}\int_{0}^{t}d\tau\left[\left(\Psi_{+}(t-\tau),\Psi^{*}_{+}(t-\tau)\right)\cdot{\bf\Phi}_{2n}^{\pm\pm}(\tau)+\left(\Psi_{-}(t-\tau),\Psi^{*}_{-}(t-\tau)\right)\cdot{\bf\Phi}_{2n+1}^{\mp\pm}(\tau)\right]$ (A9)
@@
${\bf\Phi}_{2n}^{\pm\pm}(\tau)=\int_{0}^{\tau}d\tau_{2n}e^{-i\omega_{\mp}\tau_{2n}}\psi_{\mp}(\tau_{2n})M_{\mp}(\tau_{2n})\cdots\int_{0}^{\tau}d\tau_{1}e^{-i\omega_{\pm}\tau_{1}}\psi_{\pm}(\tau_{1})M_{\pm}(\tau_{1}){\bf A}^{\pm}_{0}$ (A10)
is the convolution of $2n$ functions $e^{-i\omega_{\pm}\tau}\psi_{\pm}(\tau)M_{\pm}(\tau)$ in correspondence of $2n$ magnetic field switches in the interval $(0,\tau)$ starting with $\pm$ and ending with same sign $\pm$. We are therefore summing over all the possible realisations, or trajectories, of the fluctuations of the field. The sum in Eq. (A9) can be evaluated in Laplace transform since, using the properties of convolution, from (A10) it follows:
$\hat{\Phi}_{2n}^{\pm\pm}(s)=\left(M_{\psi_{\mp}}(s+i\omega_{\mp})\right)^{n}\left(M_{\psi_{\pm}}(s+i\omega_{\pm})\right)^{n}{\bf A}^{\pm}_{0}\;\;,\;\;\;\hat{\Phi}_{2n+1}^{\pm\pm}(s)=M_{\psi_{\pm}}(s+i\omega_{\pm})\hat{\Phi}_{2n}^{\pm\pm}(s)$ (A11)
where $M_{\psi_{\pm}}(s)=\mathcal{L}_{s}[\psi_{\pm}(\tau)M_{\pm}(\tau)]$. This leads to a geometric series in (A9) in Laplace domain that can be evaluated exactly, from the Laplace Transform of (A9):
@@
$\langle\hat{z}(s)\rangle=\frac{\langle\hat{z}^{+}(s)\rangle+\langle\hat{z}^{-}(s)\rangle}{2}$ (A13)
@@
$\langle\hat{z}(s)\rangle=\frac{\omega_{1}(s+\gamma/2)}{s^{4}+2s^{3}(i\omega_{0}+\gamma)+s^{2}(\omega_{1}^{2}-\omega_{0}^{2}+3\omega_{1}\gamma+\gamma^{2})+s\gamma(\omega_{1}^{2}-\omega_{0}^{2}+i\omega_{0}\gamma)+\omega_{1}^{2}\gamma^{2}/4}$ (A14)
@@
$\langle z(t)\rangle=\omega_{1}\sum_{i=1}^{4}\frac{e^{\lambda_{i}t}(2\lambda_{i}+\gamma)}{(\lambda_{i}-\lambda_{j\neq i})(\lambda_{i}-\lambda_{k\neq j\neq i})(\lambda_{i}-\lambda_{m\neq k\neq j\neq i})}$ (A15)
@@
$x^{4}+2(\gamma+i\omega_{0})x^{3}+(\omega_{1}^{2}-\omega_{0}^{2}+3i\omega_{0}\gamma+\gamma^{2})x^{2}+\gamma(\omega_{1}^{2}-\omega_{0}^{2}+i\omega_{0}\gamma)x+\omega_{1}^{2}\gamma^{2}/4=0$ (A16)
@@
$\langle\hat{z}(s)\rangle=\frac{i(s+\gamma)(s+i\omega_{1}+\gamma)}{(s^{4}+2s^{3}(i\omega_{0}+\gamma)+s^{2}(\omega_{1}^{2}-\omega_{0}^{2}+3\omega_{0}\gamma+\gamma^{2})+s\gamma(\omega_{1}^{2}-\omega_{0}^{2}+i\omega_{0}\gamma)+\omega_{1}^{2}\gamma^{2}/4)}$ (A17)
@@
$\langle z(t)\rangle=i\sum_{i=1}^{4}\frac{e^{\lambda_{i}t}(\lambda_{i}^{2}+\lambda_{i}(2i\gamma-\omega_{0})+i\gamma^{2}-\omega_{0}\gamma)}{(\lambda_{i}-\lambda_{j\neq i})(\lambda_{i}-\lambda_{k\neq j\neq i})(\lambda_{i}-\lambda_{m\neq k\neq j\neq i})}$ (A18)
with $\lambda_{i}$ the four roots of the same equation (A16).
VI.0.2 Evaluation of $\langle r^{2}(t)\rangle$: the case of non-ergodic and ergodic power law fluctuations